OpenStudy (anonymous):

what are the terms of sine and cosine if cos^2 theta(1+tan theta)

OpenStudy (anonymous):

\[\cos^2\theta(1+\tan^2 \theta)\]

OpenStudy (solomonzelman):

re-write tan²x as `sin²x/cos²x`

OpenStudy (anonymous):

sec^2 (x) = (1+tan^2(x)) = 1/cos^2(x) so they cancel and the answer is 1

OpenStudy (solomonzelman):

and then multiply. Then apply, sin²x+cos²x=1